OpenStudy (anonymous):

Iam Goodman's sister and need help Mrs.Williams is deciding between two field trips for her class The Science Center charges $250 plus $4 per student. The Dino Discovery Museum simply charges $9 per student. For how many students will the Science Center charge less than the Dino Discovery Museum?

OpenStudy (karatechopper):

well it depends on how many peeepl r going

OpenStudy (karatechopper):

o wait nvm

OpenStudy (lilg132):

lol that is what the question is asking karatechopper

OpenStudy (anonymous):

tats all the info it gives in the problem

OpenStudy (karatechopper):

$255 is the deiiference?

OpenStudy (anonymous):

ya mean 250?

OpenStudy (lilg132):

you mean 245 is the difference lol

OpenStudy (lilg132):

i think 28 let me chekc

OpenStudy (anonymous):

im sorry..im kinda confused..cood u explain please?

OpenStudy (lilg132):

well first these are the equations SC = 250 + 4(x) DM = 9(x)

OpenStudy (anonymous):

yes..i got that far

OpenStudy (anonymous):

huh? can u please explain?

OpenStudy (lilg132):

ok so 250 + 4(x) = 9(x) 250 + 4x = 9x 250 = 9x - 4x 250 = 5x 50 = x so to check we use the formula we got before Sc = 250 + 4(x) Dm = 9(x) Sc = 250 + 4(50) Sc = 250 + 200 Sc = 450 Dm = 9(50) Dm = 450 so the answer would be anything above 50 students the science center will charge less and just to check use 51 students in the equation Sc = 250 + 4(51) Sc =250 + 204 Sc = 454 Dm = 9(51) Dm = 459 this shows after 50 students Science center will charge less dont forget to click good answer :)

OpenStudy (anonymous):

LOLS..thanks and no problem :D